OpenStudy (anonymous):

If you know logs well can you please explain something to me? Here is the problem: rewrite and simplify the log expression: ln(5e^6) I have the answer which is 6+ln5 my question is how? I am just trying to understand how that problem gives that answer? when I look at the log properties I though it would be the power property but when i tried to solve it as such it didn't produce the same answer. can someone help me be less confused?

OpenStudy (anonymous):

\[\ln(5e^x)=\ln(5)+\ln(e^6)\] and \(\ln(e^6)=6\)

OpenStudy (anonymous):

I don't understand why it would be the product property that is used and not the power property because it has a power in the problem.

OpenStudy (anonymous):

there are two properties used here

OpenStudy (anonymous):

one is \[\log(AB)=\log(A)+\log(B)\] so \[\ln(5e^6)=\ln(5)+\ln(e^6)\]

OpenStudy (anonymous):

the second property is that the natural log, log base e, is written as \(\ln(x)\) rather than \(\log_e(x)\) so you have \(\log_e(e^x)\) which is pretty clearly 6 the natural log is the inverse of the exponential function, so it is always true that \[\ln(e^x)=x\]

OpenStudy (anonymous):

thank you so much! I couldn't figure out how the book was getting the answer

OpenStudy (anonymous):

as inverse functions, you always have \[\ln(e^x)=e^{\ln(x)}=x\] yw

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

yw again